OpenStudy (anonymous):

At Trent's garage, 3 cases of brake fluid last 4 weeks. Trent has 5 cases. How much more brake fluid should Trent order so that he has enough for 13 weeks? Round your answer up to the nearest whole number

OpenStudy (anonymous):

well we find how much 1 case lasts for by doing 4/3 = 1.33 weeks

OpenStudy (anonymous):

sorry i read that wrong, 5 * 1.33 = 6.65, 13 - 6.65 = 6.35. He needs gas for 6.35 more weeks so, 6.35/1.33 = 4.77, and we round to 5

OpenStudy (anonymous):

so we assume he will need 4.77, or 5 more cases after we round. Hope this helps :)

OpenStudy (tkhunny):

\(x\) is the number of cases to order. \(\dfrac{3\;cases}{4\;weeks} = \dfrac{x+5\;cases}{13\;weeks}\) It's a proportion!!